Calculate the values of marginal propensity to save (MPS) and marginal propensity to consume (MPC) when Y=40.

a) If the saving function is given by S° = 0.03Y^2 - 2Y + 150


Calculate the values of marginal propensity to save (MPS)

 and marginal propensity to consume (MPC) when Y=40.
Answer #1

a) Marginal propensity to save (MPS)

The marginal propensity to save (MPS) function is expressed as the derivative of the savings (S) function concerning disposable in

come (Y).

$$ M P S=\frac{d S}{d Y}=0.06 Y-2 $$

When \(Y=40\)

$$ \begin{aligned} M P S=& 0.06(40)-2=0.4 \\ 1=& M P C+M P S \\ M P C=1-& M P S=1-(0.06 Y-2) \\ &=3-0.06 Y \end{aligned} $$

When \(Y=40\)

$$ M P C=1-0.4=0.6 $$

When \(Y=40, M P S=0.4, M P C=0.6\)


answered by: aex